how d i sovle this problem 7/3=52/n n=

Answer by jim_thompson5910(35256)

7/3=52/n

7n=52*3

7n = 156

n = 156/7

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim